5.00 mL of a 0.5M HCl is added to 150.0 mL of water. What is the pH after the acid is added?
Answer #1

Final Volume of solution(inL) = 0.150+ 0.005 = 0.155L

Moles of HCl added = Molarity * Volume of HCl added(inL)

= 0.5*0.005 = 0.0025 moles HCl

So, concentration of H+ ions = moles/Total volume = 0.0025/0.155 = 0.01613 = [H+]

pH = -log[H+]

= -log(0.01613) = 1.79 (Answer)